Calculate the output of the reaction N2+3H2=2NH3 by nitrogen as a base. General pressure is 50·105 pa, equilibrium constant is K=1,64·10-4
1
Expert's answer
2019-12-12T05:46:11-0500

Formula for calculation: K=p2NH3/pN2+p3H2

Let's make a system of equations:

4p+x=50•105, where x is pressure of NH3

x2/p+27p3=1,64•10-4

After math transformations:

(50•105-4p)2=1,64 • 10-4(p+27p3)

Calculating p.

p=162297, and back to the original condition: 50•105=х+4•162297

Hence x equals x=43,5•105 Pa

So, the yield of the reaction: 43,5/50 •100% = 87%